OpenStudy (anonymous):

Two trapezoids have areas of 432 cm2 and 48 cm2. Find the similarity ratio of the sides of the trapezoids.

Directrix (directrix):

If two polygons are similar, the square of the scale factor of the two polygons is equal to the ratio of the areas of the solids. 432/48 = (scale factor)^2 (27*16) / (3*16) = (scale factor)^2 27/16 = (scale factor)^2 Take square roots of both sides to get the scale factor. @melito17